# RefineStat: Efficient Exploration for Probabilistic Program Synthesis

### v1

@@ Abstract
Probabilistic programming offers a powerful framework for modeling uncertainty, yet statistical model discovery in this domain entails navigating an immense search space under strict domain-specific constraints. When small language models are tasked with generating probabilistic programs, they frequently produce outputs that suffer from both syntactic, and semantic errors, such as flawed inference constructs. Motivated by probabilistic programmers’ domain expertise and debugging strategies, we introduce RefineStat, a language model–driven framework that enforces semantic constraints ensuring synthesized programs contain valid distributions, well-formed parameters, and then applies diagnostic-aware refinement by resampling prior or likelihood components whenever reliability checks fail. We evaluate RefineStat on multiple probabilistic-programming code-generation tasks using smaller language models (SLMs) and find that it produces programs that are both syntactically sound and statistically reliable, often matching or surpassing those from closed-source large language models (e.g., OpenAI o3).
@@ 1 Introduction
Scientific discovery often depends on our ability to express complex systems as statistical models. Finding appropriate models that are both interpretable and computationally efficient is challenging. The vision of automating model discovery has a long-standing history. Systems developed in the past have demonstrated success across various domains, such as identifying physical laws Bongard and Lipson (2007); McKinney et al. (2006); Linka et al. (2023), recovering the structure of nonlinear dynamical systems Schmidt and Lipson (2009), performing structure-aware nonparametric regression Duvenaud et al. (2013), and tackling unsupervised learning problems Grosse (2014). Despite these achievements, they typically relied on significant manual effort—experts were required to define a domain-specific language (DSL) for representing models and to engineer custom search algorithms for exploring compositions within that DSL.
@@
These limitations motivate Small Language Models (SLMs) as a practical alternative. Despite recent gains on coding tasks, they still produce semantic bugs code that runs but violates the intended statistical meaning. For example, in PyMC Salvatier et al. (2016), an SLM can generate code that places variance where a standard deviation is expected – pm.Normal(..., sigma=sigma**2) instead of pm.Normal(..., sigma=sigma). This change encodes the wrong statistical model, often inflating uncertainty and even triggering errors such as SamplingError. These issues motivate our constraints and Bayesian-workflow checks to detect and correct such errors.
Our Work: RefineStat We present RefineStat, a novel probabilistic programming synthesis framework that efficiently guides a LLM to generate probabilistic programs. RefineStat is the first to demonstrate that open-weight SLMs can synthesize reliable probabilistic programs in the Bayesian-workflow sense i.e., they satisfy standard checks. The checks include low split-$\widehat{R}$ and adequate bulk/tail ESS (good MCMC mixing), low NUTS divergences (well-behaved geometry), and strong out-of-sample fit via PSIS–LOO. These diagnostics plug into the loop of the Bayesian workflow (Gelman et al., 2020), and we adopt the improved $\widehat{R}$/ESS definitions (Vehtari et al., 2021) and PSIS–LOO for predictive assessment (Vehtari et al., 2017). Section 2 gives relevant definitions.
RefineStat produces reliable statistical programs through a two-phase approach: (1) semantically constrained generation and (2) diagnostic-aware refinement (Section 3). Our semantic constraining ensures that synthesized probabilistic programs contain valid distributions with well-formed parameters and proper variable dependencies. The diagnostic-aware refinement systematically resamples prior specifications or likelihood models when generated programs fail to meet established reliability criteria within the Bayesian workflow, thereby ensuring efficient search of probabilistic models using small langauge models.
@@ 2 Background
2.1 Language Models
Current autoregressive language models (LMs) operate on a vocabulary $V\subseteq\Sigma^{*}$ of tokens. A tokenizer converts an input prompt $O_{0}\in\Sigma^{*}$ into a sequence of tokens $t_{1},t_{2},\dots,t_{k}$. The LM $M:V^{*}\to\mathbb{R}^{|V|}$ takes this sequence and outputs scores $\mathcal{S}$ over the vocabulary: $\mathcal{S}=M(t_{1},t_{2},\dots,t_{k})$. A softmax function transforms these scores into a probability distribution, from which $t_{k+1}$ is sampled.
Bayesian Workflow A robust Bayesian analysis follows an iterative workflow of model formulation, posterior inference, diagnostic checking, and model comparison (Gelman et al., 2020). At its core, a generative program often follows a $\mathcal{D}\|\mathcal{P}\|\mathcal{L}$ structure—first fixing the observed data ($\mathcal{D}$), then sampling latent variables under the prior ($\mathcal{P}:p(z)$), and finally conditioning on the data via the likelihood ($\mathcal{L}:p(x\mid z)$). Modern probabilistic programming languages such as Stan (Carpenter et al., 2017b) and PyMC (Salvatier et al., 2016) streamline this cycle by automating MCMC sampling and offering integrated diagnostics to assess prior implications, convergence (e.g., split-$\widehat{R}$, effective sample size, BFMI, divergent NUTS transitions) (Vehtari et al., 2021; Hoffman et al., 2014; Betancourt, 2017), and predictive accuracy via PSIS-LOO (Vehtari et al., 2017). These checks guard against mis-specification and poor sampling before models are deployed for inference or decision-making.
2.2 Probabilistic Programming
Statistical modeling aims to describe relationships between variables in data through joint probability distributions that capture both observed phenomena and underlying latent structure. In probabilistic modeling, we formalize this as a joint distribution $p(x,z|\eta)$, where $x=x_{1:n}$ represents $n$ observed data points, $z=z_{1:m}$ denotes $m$ latent variables, and $\eta$ corresponds to fixed model parameters. The inferential goal is to compute the posterior distribution $p(z|x)$, which quantifies uncertainty in the latent variables conditional on observed data. Probabilistic programming languages (PPL) provide a flexible computational substrate for specifying joint distributions $p(x,z\mid\eta)$ as programs while leveraging automated inference methods (e.g., MCMC, variational inference) to compute the posterior $p(z\mid x)$ van de Meent et al. (2021). They can be categorized as internal Domain-Specific Languages (DSLs), which embed within a host language and reuse its syntax and tooling (e.g., PyMC3 Salvatier et al. (2016), NumPyro Phan et al. (2019), Pyro Bingham et al. (2019)), or as external DSLs that define their own syntax and compiler (e.g., Stan Carpenter et al. (2017b)).
This representation is particularly valuable because it enables automated model specification while leveraging established inference algorithms Gordon et al. (2014).
Probabilistic Programming Metrics We briefly define the standard probabilistic programming diagnostics and metrics used in our framework. Detailed formal definitions are provided in Appendix A.
@@ 2.2 Probabilistic Programming
    $\mathrm{BFMI}(M)$: Bayesian Fraction of Missing Information for model $M$, assessing energy transition efficiency in Hamiltonian Monte Carlo (HMC) Neal et al. (2011). HMC is a gradient-based MCMC algorithm that leverages Hamiltonian dynamics to propose coherent moves in parameter space, reducing random-walk behavior.
@@
    $\widehat{k}_{i}(M)$: Pareto shape parameter for observation $i$ in PSIS-LOO (Pareto-smoothed importance sampling leave-one-out cross-validation), quantifying reliability of importance sampling estimates. PSIS-LOO approximates exact LOO predictive densities by smoothing raw importance weights with a generalized Pareto fit to stabilize high-variance weights.
@@
Definition 1 (PSIS-LOO)
Let $\mathcal{D}=\{y_{i}\}_{i=1}^{n}$ be the observed data and $M$ a model. Draw $\{\theta^{(s)}\}_{s=1}^{S}\sim p(\theta\mid\mathcal{D})$, compute raw weights $w_{i}^{(s)}=1/p(y_{i}\mid\theta^{(s)})$, and let $\tilde{w}_{i}^{(s)}$ be the Pareto-smoothed weights (Vehtari et al., 2017). Then the PSIS-LOO estimate is
$\widehat{\mathrm{elpd}}_{\mathrm{PSIS-LOO}}\;=\;\sum_{i=1}^{n}\log\!\Bigl{[}\tfrac{1}{S}\sum_{s=1}^{S}\tilde{w}_{i}^{(s)}\,p(y_{i}\mid\theta^{(s)})\Bigr{]}.$
Instead of refitting the model $n$ times (cost $O(n)$), PSIS-LOO only needs a single full-data fit plus simple weight calculations. The fitted Pareto shape parameters $k_{i}$ are used to check reliability of this method of ELPD computation, and the common rule is that if 20% of the $k_{i}$ exceed $0.7$, the estimate should be considered unreliable.
@@ 3 RefineStat
Figure 1: The workflow of RefineStat. (1) Data and prompt are provided to the language model, which generates a probabilistic program. (2) Constrained semantic decoding enforces syntactic and semantic validity of the generated program. (3) A Bayesian reliability check diagnoses convergence, divergences, and predictive validity. If failures are detected, the model is refined by backtracking and resampling priors or likelihoods. (4) Upon passing checks, the final reliable probabilistic program is obtained.
We present RefineStat’s two main ideas. First, we prune the possible search space of probabilistic programs by enforcing semantic validity during generation, mapping validation rules to nodes in the partial parse tree and resampling problematic program fragments when constraints are violated. Second, we implement diagnostic-aware refinement, systematically resampling components of statistically unsound models to satisfy established Bayesian workflow guidelines. This integrated approach aims to achieve semantic correctness, statistical reliability, and strong predictive performance.
3.1 Problem Statement
Let $\mathcal{D}$ denote the dataset for a given statistical modeling task. The objective of RefineStat is to construct a statistical model $M$ in a probabilistic programming language that provides accurate predictive performance while quantifying uncertainty in a fully Bayesian manner.
Our approach to finding a model $M$ that explains the data will follow the standard Bayesian workflow (Gelman et al., 2020) [^1] The Bayesian workflow (Gelman et al., 2020) consists of several stages: (1) Specify model by defining likelihood and priors (in our case using an LLM). (2) Perform posterior inference, (3) Perform Posterior predictive checks. (4) Perform convergence and sampler diagnostics. (5) Compute $\widehat{\mathrm{elpd}}$(M) only if diagnostics pass. (6) Rank models by $\widehat{\mathrm{elpd}}$ (with uncertainty). The key challenge to finding such a model $M$ is to automatically compare various candidate models that an LLM produces. To automate this task, we will use a battery of diagnostics from statistical literature (Section 2.2), computed during the posterior inference in the standard Bayesian workflow (Gelman et al., 2020).
@@ 3.1 Problem Statement
Let $\mathcal{M}$ be the set of candidate models, and fix thresholds $\alpha_{R},\beta_{\mathrm{bulk}},\beta_{\mathrm{tail}},\gamma,L_{\textit{cd}}{},\epsilon$. For each $M\in\mathcal{M}$, define seven binary indicators $s_{j}(M)\in\{0,1\}$ corresponding to (see Section 2.2 for the metrics definitions):
$1.\;\max_{\phi}\widehat{R}_{\phi}(M)\leq\alpha_{R},$ $2.\;\mathrm{BFMI}(M)>\gamma,$ $3.\;\min_{\phi}\mathrm{ESS}_{\mathrm{bulk},\phi}(M)\geq\beta_{\mathrm{bulk}}$
$4.\;\mathrm{divergences}(M)=0.,$ $5.\;\min_{\phi}\mathrm{ESS}_{\mathrm{tail},\phi}(M)\geq\beta_{\mathrm{tail}},$ $6.\;\widehat{\mathrm{elpd}}(M)\text{ is finite},$
$7.\;\tfrac{1}{n}\sum_{i=1}^{n}\mathbb{I}[\widehat{k}_{i}(M)\leq L_{\textit{cd}}{}]\geq 1-\epsilon,$
Then the reliability score is $\mathcal{B}(M)=\sum_{j=1}^{7}s_{j}(M)\;\in\{0,1,\dots,7\}.$
These diagnostics can be extracted directly from the MCMC engines (e.g. Stan Carpenter et al. (2017a), PyMC Salvatier et al. (2016)). Although $\widehat{\mathrm{elpd}}$ provides a principled Bayesian measure of out-of-sample predictive accuracy, its Monte Carlo estimate can be unreliable if the sampler has not fully converged or if the importance weights are unstable (Gelman et al., 1995). To mitigate these risks, we consider elpd estimates for models that satisfy standard convergence thresholds, thus ensuring that predictive comparisons rest on reliable posterior samples.
We require each model to pass most of these seven checks: if any check fails, the corresponding $s_{j}(M)$ is zero, and the total score reflects how many diagnostics remain satisfactory. In particular, if $\widehat{\mathrm{elpd}}$(M) cannot be computed (or is infinite), then $s_{\mathrm{ELPD}}(M)=0$. A higher score means that more diagnostics have been verified, and the resulting $\widehat{\mathrm{elpd}}$ estimate can be trusted with greater confidence. We consider a model sufficiently reliable once its score exceeds a cutoff $\zeta$. We will use $\zeta=5$, to accommodate for marginal diagnostic failures while maintaining confidence in $\widehat{\mathrm{elpd}}$.
@@
Let $\mathcal{M}$ be the set of candidate models, then
@@
We finally define our objective to identify the model that attains the highest ELPD-LOO estimate within the space of models $\mathcal{M}_{\mathrm{valid}}$:

### v2

@@ Abstract
Probabilistic programming offers a powerful framework for modeling uncertainty, yet statistical model discovery in this domain entails navigating an immense search space under strict domain‐specific constraints. When small language models are tasked with generating probabilistic programs, they frequently produce outputs that suffer from both syntactic, and semantic errors, such as flawed inference constructs. Motivated by probabilistic programmers’ domain expertise and debugging strategies, we introduce RefineStat, a language model–driven framework that enforces semantic constraints ensuring synthesized programs contain valid distributions, well‐formed parameters, and then applies diagnostic‐aware refinement by resampling prior or likelihood components whenever reliability checks fail. We evaluate RefineStat on multiple probabilistic-programming code-generation tasks using smaller language models (SLMs) and find that it produces programs that are both syntactically sound and statistically reliable, often matching or surpassing those from closed-source large language models (e.g., OpenAI o3).
@@ 1 Introduction
Scientific discovery often requires expressing complex systems as statistical models. Finding appropriate models that are both interpretable and computationally efficient is challenging. The vision of automating model discovery has a long-standing history. Past approaches have demonstrated success across various domains, such as identifying physical laws Bongard and Lipson (2007); McKinney et al. (2006); Linka et al. (2023), recovering the structure of nonlinear dynamical systems Schmidt and Lipson (2009), performing structure-aware nonparametric regression Duvenaud et al. (2013), and tackling unsupervised learning problems Grosse (2014). However, they typically relied on significant manual effort – experts were required to define a domain-specific language (DSL) for representing models and engineer custom search algorithms for exploring compositions within that DSL.
@@
These limitations motivate Small Language Models (SLMs) as a practical alternative. Despite recent gains on coding tasks, they still produce semantic bugs code that runs but violates the intended statistical meaning. For example, in PyMC Salvatier et al. (2016), an SLM can generate code that places variance where a standard deviation is expected – pm.Normal(..., sigma=sigma**2) instead of pm.Normal(..., sigma=sigma). This change encodes the wrong statistical model, often inflating uncertainty and even triggering errors such as SamplingError. In addition, SLMs may produce other semantic mistakes, such as using an invalid argument name sd in place of sigma, which raises a TypeError at model construction time. These cases illustrate the need for our constraints and Bayesian-workflow checks (Gelman et al., 2020) to ensure correctness.
Our Work: RefineStat We present RefineStat, a novel probabilistic programming synthesis framework that efficiently guides a language model to generate probabilistic programs. RefineStat is the first to demonstrate that open-weight SLMs can synthesize reliable probabilistic programs in the Bayesian-workflow sense i.e., they satisfy standard checks, such as adequate effective sample size, low number of MCMC divergences and strong out-of-sample fit (Section 2 presents the full list).
RefineStat produces reliable statistical programs through a two-phase approach: (1) semantically constrained generation and (2) diagnostic-aware refinement (Section 3). In this context, semantically constrained denotes adherence to programming-language semantics (e.g., distribution validity, parameter consistency, proper data types), rather than the linguistic semantics of natural languages. Our semantic constraining ensures that synthesized probabilistic programs contain valid distributions with well-formed parameters, proper variable dependencies, and adherence to PyMC semantics. The diagnostic-aware refinement systematically resamples prior specifications or likelihood models when generated programs fail to meet established reliability criteria within the Bayesian workflow, thereby ensuring efficient search of probabilistic models using small language models.
@@ 2 Background
Language Models. Current autoregressive language models (LMs) operate on a vocabulary $V\subseteq\Sigma^{*}$ of tokens. A tokenizer converts an input prompt $O_{0}\in\Sigma^{*}$ into a sequence of tokens $t_{1},t_{2},\dots,t_{k}$. The LM $M:V^{*}\to\mathbb{R}^{|V|}$ takes this sequence and outputs scores $\mathcal{S}$ over the vocabulary: $\mathcal{S}=M(t_{1},t_{2},\dots,t_{k})$. A softmax function transforms these scores into a probability distribution, from which $t_{k+1}$ is sampled. Appendix B.1 has more details on decoding and grammar-guided generation.
Bayesian Workflow. A robust Bayesian analysis follows an iterative workflow of model specification, posterior inference, diagnostic checking, and model comparison (Gelman et al., 2020). This process can be summarized as: (1) specify the model (likelihood and priors, in our case using an LLM), (2) perform posterior inference, (3) conduct posterior predictive checks and convergence diagnostics, (4) if diagnostics pass, estimate out-of-sample fit (i.e., how well the model would predict data not used in fitting), and (5) compare and rank models by their relative out-of-sample performance (with uncertainty). Further details about diagnostics and predictive evaluation in Appendix B.2.
Probabilistic Programming. Statistical modeling aims to describe relationships between variables in data through joint probability distributions that capture both observed phenomena and underlying latent structure. In probabilistic modeling, we formalize this as a joint distribution $p(x,z|\eta)$, where $x=x_{1:n}$ represents $n$ observed data points, $z=z_{1:m}$ denotes $m$ latent variables, and $\eta$ corresponds to fixed model parameters. The inferential goal is to compute the posterior distribution $p(z|x)$, which quantifies uncertainty in the latent variables conditional on observed data. Probabilistic programming languages (PPL) provide a flexible computational substrate for specifying joint distributions $p(x,z\mid\eta)$ as programs while leveraging automated inference methods (e.g., MCMC, variational inference) to compute the posterior $p(z\mid x)$ van de Meent et al. (2021). Further details in Appendix B.3.
Probabilistic Programming Diagnostics We briefly define the standard probabilistic programming diagnostics and metrics used in our framework. Detailed formal definitions are in Appendix C:
@@
    $\mathrm{BFMI}(M)$: Bayesian Fraction of Missing Information for model $M$, assessing energy transition efficiency in Hamiltonian Monte Carlo (HMC) Neal and others (2011) algorithm.
@@
    $\widehat{k}_{i}(M)$: Pareto shape parameter for observation $i$ in PSIS-LOO (Pareto-smoothed importance sampling leave-one-out cross-validation), quantifying reliability of importance sampling estimates. PSIS-LOO approximates exact LOO predictive densities by smoothing raw importance weights with a generalized Pareto fit to stabilize high-variance weights (Definition 1 below).
@@
Definition 1 (PSIS-LOO (Vehtari et al., 2017))
Let $\mathcal{D}=\{y_{i}\}_{i=1}^{n}$ be the observed data and $M$ a model. Draw $\{\theta^{(s)}\}_{s=1}^{S}$ from $p(\theta\mid\mathcal{D})$, compute raw importance weights $w_{i}^{(s)}=1/p(y_{i}\mid\theta^{(s)})$, and let $\tilde{w}_{i}^{(s)}$ denote the Pareto-smoothed weights, obtained by replacing the largest tail weights with a generalized Pareto fit. Then the PSIS-LOO estimate is $\widehat{\mathrm{elpd}}_{\mathrm{PSIS\text{-}LOO}}\;=\;\sum_{i=1}^{n}\log\!\left[\frac{1}{S}\sum_{s=1}^{S}\tilde{w}_{i}^{(s)}\,p(y_{i}\mid\theta^{(s)})\right].$
Instead of refitting the model $n$ times, PSIS-LOO relies on a single full-data fit and stabilized importance weights. The generalized Pareto fit yields shape parameters $k_{i}$, which assess the reliability of the approximation; following standard guidance, the estimate is considered unreliable if about 20% of the $k_{i}$ exceed $0.7$.
@@ 3 RefineStat
Figure 1: RefineStat workflow: (1) A user provides data and prompt to the language model, which generates a probabilistic program. (2) Constrained semantic decoding enforces syntactic and semantic validity of the generated program. (3) A Bayesian reliability check diagnoses convergence, divergences, and predictive validity. If failures are detected, the model is refined by backtracking and resampling priors or likelihoods. (4) Upon passing checks, we get final reliable probabilistic program.
Figure 1 presents RefineStat’s two main ideas. First, we prune the search space of possible probabilistic programs by enforcing semantic validity during generation, mapping validation rules to nodes in the partial parse tree and resampling problematic program fragments when constraints are violated. Second, we implement diagnostic-aware refinement, systematically resampling components of statistically unsound models to satisfy Bayesian Workflow guidelines. This integrated approach aims to improve semantic correctness, statistical reliability, and yield strong predictive performance. An illustrative example is provided in Appendix D.1.
Problem Statement. Let $\mathcal{D}$ denote the dataset for a given statistical modeling task. The objective of RefineStat is to construct a statistical model $M$ in a probabilistic programming language that provides accurate predictive performance while quantifying uncertainty in a fully Bayesian manner.
Our approach to finding a model $M$ that explains the data will follow the standard Bayesian workflow (Gelman et al., 2020) The key challenge to finding such a model $M$ is to automatically compare various candidate models that an LLM produces. To automate this task, we will use a battery of diagnostics from statistical literature (Section 2), computed during the posterior inference in the standard Bayesian workflow (Gelman et al., 2020).
@@
Let $\mathcal{M}$ be the set of candidate models, and fix thresholds $\alpha_{R},\beta_{\mathrm{bulk}},\beta_{\mathrm{tail}},\gamma,L_{\textit{cd}}{},\epsilon$. For each $M\in\mathcal{M}$, we define seven indicator functions $s_{j}(M)\in\{0,1\}$ by $\mathbb{I}[A]=\begin{cases}1,&\text{if event }A\text{ holds},\\ 0,&\text{otherwise},\end{cases}$ and set
$1.\;\;s_{1}(M)$ $=\mathbb{I}\!\left[\max_{\phi}\widehat{R}_{\phi}(M)\leq\alpha_{R}\right],$ $2.\;\;s_{2}(M)$ $=\mathbb{I}\!\left[\mathrm{BFMI}(M)>\gamma\right],$
$3.\;\;s_{3}(M)$ $=\mathbb{I}\!\left[\min_{\phi}\mathrm{ESS}_{\mathrm{bulk},\phi}(M)\geq\beta_{\mathrm{bulk}}\right],$ $4.\;\;s_{4}(M)$ $=\mathbb{I}\!\left[\mathrm{divergences}(M)=0\right],$
$5.\;\;s_{5}(M)$ $=\mathbb{I}\!\left[\min_{\phi}\mathrm{ESS}_{\mathrm{tail},\phi}(M)\geq\beta_{\mathrm{tail}}\right],$ $6.\;\;s_{6}(M)$ $=\mathbb{I}\!\left[\tfrac{1}{n}\!\sum_{i=1}^{n}\mathbb{I}[\widehat{k}_{i}(M)\leq L_{\textit{cd}}{}]\geq 1-\epsilon\right],$
$7.\;\;s_{7}(M)$ $=\mathbb{I}\!\left[\widehat{\mathrm{elpd}}(M)\text{ is finite}\right]$ Then the reliability score is $\mathcal{B}(M)=\sum_{j=1}^{7}s_{j}(M)\$.
These diagnostics can be extracted directly from the MCMC engines (e.g. Stan Carpenter et al. (2017b), PyMC Salvatier et al. (2016)). Although $\widehat{\mathrm{elpd}}$ provides a principled Bayesian measure of out‑of‑sample predictive accuracy, its Monte Carlo estimate can be unreliable if the sampler has not fully converged or if the importance weights are unstable (Gelman et al., 1995). To mitigate these risks, we consider elpd estimates for models that satisfy standard convergence thresholds, thus ensuring that predictive comparisons rest on reliable posterior samples.
We require each model to pass most of these seven checks: if any check fails, the corresponding $s_{j}(M)$ is zero, and the total score reflects how many diagnostics remain satisfactory. The final check concerns the availability of the $\widehat{\mathrm{elpd}}(M)$ estimate; if $\widehat{\mathrm{elpd}}(M)$ cannot be computed or is infinite, then $s_{\mathrm{ELPD}}(M)=0$, and the model is treated as failing that diagnostic. A higher overall score indicates that more diagnostics have passed, so when $\widehat{\mathrm{elpd}}(M)$ is available, the resulting estimate can be trusted with greater confidence. We consider a model reliable once its score exceeds a cutoff $\zeta$. We use $\zeta=5$ to allow marginal diagnostic failures while maintaining confidence in the reported $\widehat{\mathrm{elpd}}$.
@@
For a set of candidate models $\mathcal{M}$, a valid model space is:
@@
We finally define our objective to identify the model that attains the highest ELPD‑LOO estimate within the space of models $\mathcal{M}_{\mathrm{valid}}$: